calculate conductivity of gold at 200°c.

1
Expert's answer
2021-04-29T10:42:14-0400

The resistivity of gold at 20°C is 2.3·10-8 Ohm·m. The temperature coefficient of resistance is 0.003715/°C. Thus, the resistivity at 200°C will be


"\rho=\rho_0(1+\alpha T),"

and the conductivity


"\gamma=\frac1\rho=\frac{1}{\rho_0(1+\alpha T)},\\\space\\
\gamma=\frac{1}{2.3·10^{-8}(1+0.003715·200)}=24.94·10^6\text{ Sm/m}."